# Probing flavored regimes of leptogenesis with gravitational waves from cosmic strings

### v1

@@ Abstract
In order to robustly claim the spectral fall as a signature of $M_{N}$-dependent matter domination, we must require both $t_{\rm dec}>l_{c}/\alpha$ and $t_{\rm dec}>t_{\rm fric}$. The first condition leads to the constraint
$M_{N}\gtrsim M^{c}_{N}\left(\frac{T_{\rm dec}}{\rm GeV}\right)^{2/5}\left(\frac{g^{\prime}}{10^{-5}}\right)^{7/10}h(g^{\prime})^{-2/5}\,,$ (11)
with $M^{c}_{N}=3.8\times 10^{7}$ GeV. Similar constraints can be derived for loops containing kinks, which however are less stringent. The second condition $t_{\rm dec}>t_{\rm fric}$ implies
$M_{N}\gtrsim M^{\rm fric}_{N}\left(\frac{T_{\rm dec}}{\rm GeV}\right)^{1/2}\left(\frac{g^{\prime}}{10^{-5}}\right)h(g^{\prime})^{-1/2}\,,$ (12)
with $M^{\rm fric}_{N}=2.0\times 10^{4}$ GeV.
@@
The parameter space of our model is also constrained by the recent PTA observations of stochastic GW which disfavor cosmic strings due to a different spectral slope [31]. Thus, PTA data require $\Omega_{GW}<\Omega_{GW}^{\rm PTA}$ at $f=30$ nHz giving $G\mu\lesssim 7.0\times 10^{-11}$. This translates to the constraint on $M_{N}$ as
$M_{N}\lesssim M^{\rm PTA}_{N}\left(\frac{g^{\prime}}{10^{-5}}\right)h(g^{\prime})^{-1/2}\,,$ (13)
whit $M^{\rm PTA}_{N}\simeq 5.6\times 10^{8}$ GeV. All these constrained are satisfied by the benchmark cases shown in Fig. 1.
@@
We note that there could be additional contribution to the GW spectrum at very high frequencies due to the cosmic string loops in the first radiation epoch before $\Phi$ starts to dominate. However, the amplitude at higher frequencies is suppressed not only because of the entropy production, but also because the hard particle production cut-off makes the spectrum red.
Figure 2: The white (shaded) regions correspond to the allowed (excluded) model parameter space in the $T_{\rm dec}$-$M_{N}$ plane for $g^{\prime}=10^{-6},10^{-5},10^{-4}$ (left to right). The sensitivity of LISA [36], DECIGO [37] and ET [38] are shown with the black curves. The diagonal colored lines represent the relation Eq. (5) predicted by our model. The symbols correspond to the benchmark cases whose GW spectra are reported in Fig. 1. The black horizontal thin lines mark the separation between the two-flavor (2FL) and the three-flavor (3FL) regimes of thermal leptogenesis.
Results and discussion. In Fig. 2, we show the sensitivities of LISA, DECIGO and ET in the plane $T_{\rm dec}$-$M_{N}$, assuming $g^{\prime}=10^{-6}$ (left panel), $g^{\prime}=10^{-5}$ (middle panel) and $g^{\prime}=10^{-4}$ (right panel), together with all the conditions and constraints discussed above. We underline that the dependence of the GW sensitivities on $g^{\prime}$ comes from the relation $M_{N}=(G\mu{g^{\prime}}^{2}M_{\rm Pl}^{2}/\pi/h(g^{\prime}))^{1/2}$. The white regions in each plot can be generally regarded as allowed parameter space when taking $T_{\rm dec}$ and $M_{N}$ as independent parameters. However, in our model, they are related by Eq. (5) which restricts the allowed parameter space along the solid colored lines in the plots. Overall, the requirement of negligible particle production from cosmic string, the recent PTA data, and the BBN constraint limit the parameter space to be $g^{\prime}\sim[10^{-6},10^{-4}]$ and $M_{N}/{\rm GeV}\sim[10^{7},10^{10}]$.
We find that, irrespective of the values of $g^{\prime}$, going from light to heavy $M_{N}$ region along the colored lines, the spectral break frequency $f_{\rm dec}$ increases according to Eq.s (5) and (10). This trend allows us to point out that if the spectral break occurs in the LISA band, it most likely corresponds to $M_{N}\lesssim 10^{9}$ (3FL regime), as proven by the allowed segment of the blue line in the left plot. With a similar argument, a spectral break in the ET band would correspond to a 2FL regime of leptogenesis, as depicted by the red line in the right plot. An intermediate situation with $g^{\prime}=10^{-5}$, both the 2FL and 3FL regions would provide a spectral break in the DECIGO frequency band. Therefore, in this frequency band, the spectral break signature of 2FL and 3FL becomes ambiguous.
@@
The allowed parameter space of our model can be translated into allowed regions for the GW plateau amplitude $\Omega_{\rm GW}^{\rm plt}h^{2}$ and the break frequency $f_{\rm dec}$ (see Supplemental Material II for details). Usually, they are limited from above by the PTA data. However, in our model, they are also bounded from below by the BBN constraint due to the matter-dominated epoch and by the requirement that the spectral break is not due to the particle production from cosmic strings. We find that our scenario implies $10^{-12}\lesssim\Omega_{\rm GW}^{\rm plt}h^{2}\lesssim 10^{-10}$ and $10^{-3}\lesssim f_{\rm dec}/{\rm Hz}\lesssim 10^{3}$ which we define as the Gravitational Waves windows for Flavored Regimes of Leptogenesis (GWFRL). These ranges, within which a robust GW signature of high-scale leptogenesis is achieved, will be fully explored by planned GW experiments such as LISA, DECIGO and ET. These GWFRL windows also trivially satisfy two important constraints. First, $G\mu<6.3\times 10^{-3}g^{\prime}h(g^{\prime})$ due to the Hubble friction being always less than $m_{\Phi}$ at $T\simeq T_{c}$. Second, the ratio of the $\Phi$ vacuum energy to the radiation at $T_{c}$, which in our case is $\rho_{\Phi}(T_{c})/\rho_{R}(T_{c})=4.7\times 10^{-4}g^{\prime}$, is always smaller than one, thus avoiding a second period of inflation.
There could be possible corrections to the GWFRL windows relaxing some of our assumptions. First, we have considered $\lambda={g^{\prime}}^{3}$ but a more rigorous parameter space scan could be performed to find a smooth transition of the field and at the same time being consistent with other gauge coupling dependent constraints and a negligible $\Phi\rightarrow Z^{\prime}Z^{\prime}$ decays. Second, we have presented our results on GWs for $k=1$ mode. However, when summing over a large number of modes, the overall amplitude and the spectral break frequency would change slightly. Third, we have assumed $f_{N}=g^{\prime}$. Weakening this assumption but being still consistent with Eq. (3) would anyway not modify the overall trend of the parameter space. Indeed, smaller $f_{N}$ ($M_{N}$) would correspond to smaller $T_{\rm dec}$ and therefore smaller $f_{\rm dec}$, meaning a 3FL regime likely to show up the spectral break at lower frequencies than the 2FL regime. Summarizing, we point out that the GWFRL windows we have derived in the present paper are quite robust against our assumptions. In conclusions, we demonstrate that different flavor regimes of high-scale leptogenesis can be realistically probed with GW signatures from cosmic strings, thus establishing synergy between low-energy neutrino physics and gravitational waves.
@@
The Supplemental Material is organized as follows. In Sec. I, we discuss the evolution of the radiation and the scalar field energy densities, and the produced entropy with the inverse of temperature. We derive the analytical expressions for the decay temperature $T_{\rm dec}$ and the amount $\kappa$ of the produced entropy which match the numerical results well. In Sec. II, we detail the constraints leading to the Gravitational Waves window for Flavored Regimes of Leptogenesis (GWFRL).
@@
[…]
@@
with $a$ and $g_{*s}$ being the scale factor and the number of degree of freedom, respectively. The production of entropy from the $\Phi$ decays is computed by solving
@@
and computing the ratio of $\tilde{S}\sim a^{3}/z^{3}$ after and before the scalar field decay. The amount of entropy production $\kappa=\tilde{S}_{\rm after}/\tilde{S}_{\rm before}$ can be analytically quantified as
$\mathbb{\kappa}\simeq\left[\frac{\rho_{R}\left(T_{c}\right)}{\rho_{\Phi}\left(T_{c}\right)}\frac{T_{\rm dec}}{T_{c}}\right]^{-1}\simeq\frac{\rho_{\Phi}\left(T_{c}\right)}{3^{1/4}\left(\frac{30}{\pi^{2}g_{*}}\right)^{-3/4}T_{c}^{3}\sqrt{\Gamma_{\Phi}^{HH}\tilde{M}_{\rm Pl}}}\simeq 140\left(\frac{g^{\prime}}{10^{-3}}\right)^{-1/4}\left(\frac{M_{N}}{10^{9}~{}{\rm GeV}}\right)^{-1/2}\,,$ (S5)
by considering $\Gamma_{\Phi}^{HH}\simeq\mathcal{H}$ which gives the decay temperature $T_{\rm dec}$
$T_{\rm dec}\simeq\left(\frac{90}{\pi^{2}g_{*}}\right)^{1/4}\sqrt{\Gamma_{\Phi}^{HH}\tilde{M}_{\rm Pl}}\simeq 6.35\left(\frac{g^{\prime}}{10^{-5}}\right)^{3/4}\left(\frac{M_{N}}{10^{9}~{}{\rm GeV}}\right)^{3/2}{\rm GeV}\,,$ (S6)
@@
with $\tilde{M}_{\rm Pl}=2.4\times 10^{18}~{}{\rm GeV}$ being the reduced Planck constant, and $g_{*}\simeq 106.75$ is the effective degrees of freedom that contribute to the radiation. The evolution of the radiation, the scalar field normalized energy densities $\Omega_{i}=\rho_{i}/\rho_{\rm tot}$, and the entropy production factor $\kappa$ corresponding to the benchmarks discussed in the main text are shown in Fig. S1.
II Gravitational Waves window for Flavored Regimes of Leptogenesis (GWFRL)
@@
Figure S1: Evolution of the radiation and the scalar field normalized energy densities $\Omega_{i}=\rho_{i}/\rho_{\rm tot}$, and the entropy production factor $\kappa$ corresponding to the benchmarks discussed in the main text. Because only the ratio of entropies before and after the $\Phi$ decay matters, we set $\kappa(T_{C})=1$. The vertical lines represent the cut-off temperatures $T_{\rm fric}\gg T_{\rm fric}\gg T_{\rm dec}$, implying that the fall in the GW spectra occurs due to the $M_{N}$-dependent matter domination.
Figure S2: The white regions are allowed windows for the GW plateau amplitude (left) and the spectral break frequencies (right), referred to as Gravitational Waves Window for Flavored Regimes of Leptogenesis (GWFRL) in the main text. These windows are derived from the spectral features of GWs radiated from the fundamental mode oscillation of cosmic string loops. The symbols correspond to the benchmark cases discussed in the main text.
We here detail the upper and lower bounds on the the GW plateau amplitude $\Omega^{\rm plt}_{GW}h^{2}$ and the break frequency $f_{\rm dec}$ which allow us to define the GWFRL windows reported in Fig. S2. These regions are bounded from below by either the BBN or the particle production constraint, and from above by the PTA observations. The lower bounds on the GW amplitude (from Eq. (9)) come from the lower bound on $G\mu$ imposed by BBN and particle production from cosmic strings, which read
$(G\mu)^{\rm BBN}$ $\gtrsim$ $3.8\times 10^{-14}\left(\frac{g^{\prime}}{10^{-5}}\right)^{-3}h(g^{\prime})\,,$ (S7)
$(G\mu)^{c}$ $\gtrsim$ $6.4\times 10^{-16}\left(\frac{g^{\prime}}{10^{-5}}\right)^{3}h(g^{\prime})^{-1}\,.$ (S8)
Similarly, the lower bounds on the spectral break frequency are derived as
$f^{\rm BBN}_{\rm dec}$ $\gtrsim$ $3.2\times 10^{-2}\left(\frac{g^{\prime}}{10^{-5}}\right)^{3/2}h(g^{\prime})^{-1/2}~{}{\rm Hz}\,,$ (S9)
$f^{c}_{\rm dec}$ $\gtrsim$ $1.2\times 10^{-2}\left(\frac{g^{\prime}}{10^{-5}}\right)^{3}h(g^{\prime})^{-1}~{}{\rm Hz}\,.$ (S10)
The upper bound on the plateau amplitude is given by
$(G\mu)^{\rm PTA}\lesssim 7.0\times 10^{-11}\,.$ (S11)
while for $f^{\rm PTA}_{\rm dec}$ we have
$f^{\rm PTA}_{\rm dec}\lesssim 2.0\times 10^{-1}\left(\frac{g^{\prime}}{10^{-5}}\right)^{9/4}h(g^{\prime})^{-3/4}\rm Hz\,.$ (S12)

### v2

@@ Abstract
To robustly claim the spectral fall as a signature of $M_{N}$-dependent matter domination, we must require $t_{\rm dec}>l_{c}/\alpha$. This is equivalent to imposing a constraint on the parameter space such that $T_{\rm cusp}/T_{\rm dec}>1$, where $T_{\rm cusp}$ can be derived as
$T_{\rm cusp}\simeq 25\,{\rm GeV}\left(\frac{G\mu}{10^{-12}}\right)^{5/4}\left(\frac{g^{\prime}}{10^{-4}}\right)^{3/4}h(g^{\prime})^{-1/4}\,.$ (10)
@@
For loops containing kinks, one can obtain a similar constraint, which however is less stringent than Eq. (10). The parameter space of our model is also constrained by the recent PTA observation of stochastic GWs, which disfavors stable cosmic strings due to a different spectral slope [34]. Thus, PTA data require $\Omega_{\rm GW}\lesssim\Omega_{\rm GW}^{\rm PTA}$ at $f=30$ nHz– giving $G\mu\lesssim 7.0\times 10^{-11}$.
@@
We note that there could be an additional contribution to the GW spectrum at very high frequencies from the cosmic string loops in the first radiation epoch before $\Phi$ starts to dominate. However, the amplitude at higher frequencies is suppressed by the entropy production and the particle production cut-off.
Figure 3: The GWFRL windows for different values of the leptogenesis scale $M_{N}$, scanning over the model parameters $g^{\prime}$ and $f_{N}$. The black lines show the sensitivity of LISA, DECIGO and ET detectors (black lines). The grey shaded region is excluded by current PTA data.
@@
Results and discussion: In Fig. 2, we show the allowed parameter space for $g^{\prime}=10^{-4}$, in agreement with all the model conditions and constraints discussed above. For a given gauge coupling, the spectral break frequency $f_{\rm dec}$ and the GW plateau amplitude $\Omega_{\rm GW}^{\rm plt}$ are uniquely determined by the model parameters $M_{N}$ and $f_{N}$. Usually, they are limited from above by the PTA data. However, in our model, they are also bounded from below by the BBN constraint on the matter-dominated epoch and by the requirement that the spectral break is not due to the particle production from cosmic strings. Scanning over different values of the gauge coupling (see Supplemental Material III for the gauge coupling dependence of the parameter space), we find that our scenario implies $10^{-3}\lesssim f_{\rm dec}/{\rm Hz}\lesssim 10^{3}$ and $10^{-13}\lesssim\Omega_{\rm GW}^{\rm plt}h^{2}\lesssim 10^{-10}$, which we define as the Gravitational Waves windows for Flavored Regimes of Leptogenesis (GWFRL). These ranges, within which a robust GW signature of high-scale leptogenesis is achieved, will be fully explored by planned GW experiments such as LISA, DECIGO and ET. These GWFRL windows also trivially satisfy two important constraints. First, $G\mu<6.3\times 10^{-3}g^{\prime}h(g^{\prime})$ due to the Hubble friction being always less than $m_{\Phi}$ at $T\simeq T_{c}$. Second, the ratio of the $\Phi$ vacuum energy to the radiation at $T_{c}$; $\rho_{\Phi}(T_{c})/\rho_{R}(T_{c})=4.7\times 10^{-4}g^{\prime}$, is always smaller than one, thus avoiding a second period of inflation.
In Fig. 3, we report the allowed regions for $f_{\rm dec}$ and $\Omega_{\rm GW}^{\rm plt}$, once we fix the value of the leptogenesis scale $M_{N}$ (regions with different colors) and take the gauge coupling to vary from $10^{-7}$ to $10^{-1}$. The black lines (from left to right) represent the sensitivity of LISA, DECIGO, and ET detectors, respectively. Irrespective of the values of the gauge coupling, going from light to heavy $M_{N}$, the maximum allowed value for the spectral break frequency $f_{\rm dec}$ increases. Remarkably, this trend allows us to point out that a spectral break in the LISA (DECIGO) frequency band would be (most likely) a striking signature of the 3FL leptogenesis regime ($M_{N}\leq 10^{9}~{}{\rm GeV}$). On the other hand, at the ET frequency band, both the 3FL and 2FL regimes leave their signatures distinct in signal strength, with a smaller GW amplitude for the 3FL case. As shown with the BP1 (2FL) and BP2 (3FL) cases, despite corresponding to the same $f_{\rm dec}$, they have different plateau amplitudes (see Fig. 1). On the other hand, the 2FL regime (corresponding to $g^{\prime}\gtrsim 10^{-3}$) is highly disfavored by the PTA data, which already excluded the 1FL case.
The GWFRL windows obtained in the present analysis are robust at the leading order, nevertheless, they can be further improved in future works. We have considered $\lambda={g^{\prime}}^{3}$, but a more rigorous parameter space scan could be performed to find a smooth transition of the field while being consistent with other gauge coupling dependent constraints (see Supplemental Material III for a brief discussion). Moreover, one can account for the GW emission from the higher $k$-modes. However, we find that within an instantaneous scaling network (instantaneous change in $A_{\beta}$ during the radiation-to-matter transition), summing a large number of modes makes the PTA constraint stronger by a factor of $\sim\zeta(4/3)$ and lowers the $f_{\rm dec}$ by a factor of $\sim 2.5$. The latter has been derived by taking a $10\%$ deviation from the near scale-invariant spectrum. In addition, the spectrum falls as $f^{-1/3}$ instead of $f^{-1}$ at very high frequencies [15, 17] with no impact on our results based on the detectability of the signal at lower frequencies.
In conclusion, we have demonstrated that different flavor regimes of high-scale leptogenesis could produce GW signatures with spectral features depending on the mass acle of the heavy neutrinos. This is an interesting aspect of the simple realization of the seesaw model based on $U(1)_{B-L}$ symmetry such that different flavored regimes could be probed without invoking any flavor symmetry to reduce the number of free parameters in the seesaw Lagrangian. On the other hand, we have established synergy between primordial gravitational waves and low-energy neutrino physics in seesaw models with flavor symmetries. It is important to emphasize that the predicted GW signal should not be regarded as a definitive indicator, as similar spectral features could arise in other BSM models involving matter domination. Furthermore, in this work, we do not delve into the high-frequency behavior of the GW spectra pertaining to different flavor regimes, nor do we perform a statistical analysis to address uncertainties in reconstructing weak GW signals (e.g., BP3 in Fig. 1). These aspects will be thoroughly explored in an extended version of this work.
@@
The Supplemental Material is organized as follows. In Sec. I, we discuss the evolution of the radiation and the scalar field energy densities, and the produced entropy with the inverse of temperature. In Sec. II, we justify why the cosmic strings predicted by our model can be treated as thin strings and we detail their cosmological evolution. In Sec. III, we investigate the dependence of the Gravitational Waves windows for Flavored Regimes of Leptogenesis (GWFRL) on the gauge coupling.
@@ I Scalar field evolution and entropy production
where $\rho_{R}$ and $\rho_{\Phi}$ are the energy densities of radiation and the scalar field as a function of $z=T_{c}/T$, $\mathcal{H}$ is the Hubble parameter, $s$ is the entropy density of the thermal bath, and the temperature-time relation has been derived from the third of Eq. (S1) as
@@
with $a$ and $g_{*s}$ being the scale factor and the number of entropy degrees of freedom, respectively. The production of entropy from the $\Phi$ decays is computed by solving
@@
and computing the ratio of $\tilde{S}\sim a^{3}/z^{3}$ after and before the scalar field decay. The amount of entropy production $\kappa=\tilde{S}_{\rm after}/\tilde{S}_{\rm before}$ can be estimated as
$\kappa=\frac{T_{\rm dom}}{T_{\rm dec}}\Theta\left[T_{\rm dom}-T_{\rm dec}\right]\quad{\rm with}\quad T_{\rm dom}=\frac{\rho_{\Phi}(T_{c})}{\rho_{R}(T_{c})}T_{c}\,,$ (S5)
where $T_{c}$ is the critical temperature of the $U(1)_{B-L}$ phase transition, $T_{\rm dom}$ is the temperature at which the scalar field dominates the universe’s energy density, and the $\Theta$ function represents that the expression for $\kappa$ is valid for $T_{\rm dom}>T_{\rm dec}$ only. The temperature $T_{\rm dec}$ is the solution of the equation $\Gamma_{\Phi}^{HH}(T_{\rm dec})\simeq\mathcal{H}(T_{\rm dec})$, and takes the expression
$T_{\rm dec}=\tilde{T}_{\rm dec}~{}\mathcal{W}\left(\frac{\Lambda}{\widetilde{T}_{\rm dec}}\right)\,,$ (S6)
with
$\widetilde{T}_{\rm dec}=\left(\frac{90}{\pi^{2}g_{*}}\right)^{1/4}\left[\tilde{M}_{\rm Pl}~{}\Gamma_{0}~{}\frac{f_{N}^{6}}{\lambda}\left(\frac{v_{\Phi}}{\rm 10^{13}~{}GeV}\right)^{2}\left(\frac{m_{\Phi}}{\rm 10^{8}~{}GeV}\right)\right]^{1/2}\,,$ (S7)
@@
where $\tilde{M}_{\rm Pl}=2.4\times 10^{18}~{}{\rm GeV}$ is the reduced Planck constant, $g_{*}\simeq 106.75$ is the number of effective degrees of freedom that contribute to the radiation, and $\mathcal{W}({\Lambda}/{\widetilde{T}_{\rm dec}})$ is a Lambert function that modulates the $T_{\rm dec}$ according to the imposed renormalisation condition. The evolution of the radiation, the scalar field normalized energy densities $\Omega_{i}=\rho_{i}/\rho_{\rm tot}$, and the entropy production factor $\kappa$ corresponding to the benchmarks discussed in the main text are shown in Fig. S1.
@@
Figure S1: Evolution of the radiation and the scalar field normalized energy densities $\Omega_{i}=\rho_{i}/\rho_{\rm tot}$, and the entropy production factor $\kappa$ corresponding to the benchmarks discussed in the main text. Because only the ratio of the entropy before and after the $\Phi$ decay matters, we set $\kappa(T_{c})=1$. The vertical long-dashed, dashed and solid lines represent the cut-off temperatures $T_{\rm dom}$, $T_{\rm cusp}$ and $T_{\rm dec}$, respectively. The condition $T_{\rm cusp}\geq T_{\rm dec}$ ($z_{\rm cusp}\leq z_{\rm dec}$) implies that the fall in the GW spectra occurs due to the $M_{N}$-dependent matter domination.
We also provide an order of magnitude estimate on the required level of quasi-degeneracy in the RH neutrinos to produce the observed baryon asymmetry. The baryon-to-photon ratio in the resonance regime of leptogenesis can be written as
$\eta_{B}=10^{-2}\varepsilon_{i}\zeta_{i}\kappa^{-1}\,\,\,{\rm with}\,\,\,\varepsilon_{i}\simeq\frac{3M_{i}m_{\nu}}{8\pi v_{h}^{2}\delta}\,,$ (S8)
where $\varepsilon_{i}$ is the CP asymmetry parameter, $\zeta_{i}\simeq 10^{-1}$ is the efficiency of lepton asymmetry production that takes into account the washout processes, $\delta=(M_{j}-M_{i})/{M_{i}}$ is the level of quasi degeneracy in the RH neutrino masses, and $\kappa$ is the injected entropy by $\Phi$ decay as defined in Eq. (S5). Given the observed value $\eta_{B}\simeq 6.3\times 10^{-10}$ [2], using Eq. (S8), we obtain
$\delta\simeq 6.25\times 10^{-1}\kappa^{-1}\left(\frac{M_{N}}{10^{9}\rm GeV}\right).$ (S9)
Therefore, e.g., for BP4 with $\kappa\simeq 10^{6}$, we get $\delta\simeq 10^{-7}$ which is six orders of magnitude smaller than the standard scenarios with $\kappa=1$ (no entropy production).
II The string thickness and evolution
We name the strings produced in our model (with small $\lambda$) as “thick” strings because for a given $v_{\Phi}$ their width $\delta_{w}\sim(\sqrt{\lambda}v_{\Phi})^{-1}$ is larger than what is generally considered for $\lambda\simeq 1$. However, in an expanding Universe, a string is considered to be thin when $\delta_{w}\sim(\sqrt{\lambda}v_{\Phi})^{-1}\ll\mathcal{H}(T_{F})^{-1}$ is satisfied, where $\mathcal{H}(T_{F})^{-1}$ is the Hubble radius at the string formation temperature $T_{F}$. In our model, the above condition which translates to $G\mu\ll 10^{-3}g^{\prime}$ with $T_{F}\sim T_{c}$, is trivially satisfied for the entire allowed parameter space. Thus, while the predicted strings in our model are technically thick, they effectively behave like thin strings. We intuitively expect that the dynamics of these predicted strings will closely resemble those typically associated with thin strings. This perspective is commonly presented in the literature addressing small $\lambda$.
Let us assume that the strings are “thin” so they can be treated as the commonly-considered Nambu-Goto ones. In our scenario, the scalar field does not immediately dominate the energy density, but it takes time as defined by the temperature $T_{\rm dom}$ given in Eq. (S5). Therefore, the interval $T_{c}>T>T_{\rm dom}$ is a radiation-dominated epoch. Matter domination is obtained only within the temperature interval $T_{\rm dom}>T>T_{\rm dec}$. Typically, the string network enters the scaling regime unless the network is damped by friction defined by a temperature $T_{\rm fric}\sim\mu/M_{\rm pl}$. Hence, the network starts to emit GWs for $T_{\rm emission}\lesssim T_{\rm fric}$. In our scenario, $T_{\rm emission}$ could be either $T_{\rm emission}>T_{\rm dom}$ or $T_{\rm emission}<T_{\rm dom}$, with the preference for the former equation for majority of our parameter space. Since the evolution of the number densities of string loops from radiation to matter is well established, we can clearly evolve the network from $T>T_{\rm dom}$ to $T<T_{\rm dom}$, as the underlying process remains the same: evolving a string network from the equation of state $w=1/3$ (radiation) to $w=0$ (matter). It is crucial that $T_{\rm emission}$ remains greater than $T_{\rm dec}$. If $T_{\rm emission}<T_{\rm dec}$, the spectral break would be dictated by $T_{\rm emission}$ instead of $T_{\rm dec}$, with the latter deeply important for linking leptogenesis with the GW spectral features. In our scenario, we consistently have $T_{\rm emission}>T_{\rm dec}$ and the emitted GWs for $T>T_{\rm dec}$ are suppressed.
III Impact of the gauge coupling on the GWFRL window
In Fig. S2, we report the GWFRL windows for different choices of gauge coupling. We find that the largest allowed ranges of the spectral break frequency and the GW plateau amplitude are robust against different choices of gauge coupling. This justifies our definition of the GWFRL window in the main letter as $10^{-3}\lesssim f_{\rm dec}/{\rm Hz}\lesssim 10^{3}$ and $10^{-13}\lesssim\Omega_{\rm GW}^{\rm plt}h^{2}\lesssim 10^{-10}$. Moreover, we note that the PTA data implies that the gauge coupling is constrained to be smaller than about $3\times 10^{-3}$ (see the bottom-right panel). This further implies that the one-flavor leptogenesis regime ($M_{N}\geq 10^{12}~{}{\rm GeV}$) is already excluded by observations. A pure 3FL (2FL) regime is achieved for small (large) gauge coupling as shown in the top-left (bottom-right) panel. Both 3FL and 2FL regimes are allowed in the case of intermediate gauge coupling (top-right and bottom-left panels).
Figure S2: The allowed parameter space (GWFRL windows) in the $f_{\rm dec}$-$\Omega_{\rm GW}^{\rm plt}\,h^{2}$ plane for different values of the gauge coupling $g^{\prime}$ with the leptogenesis scale $M_{N}$ color-coded. The different black lines display the constraints according to the discussed model requirements and the PTA measurements. In the top panels, the constraint on the coupling $f_{N}$ is less restrictive being outside the displayed ranges. The orange solid line (if present) marks the transition between three-flavor and two-flavor leptogenesis regimes.
Let us also mention that the coupling hierarchy $\lambda\simeq{g^{\prime}}^{3}$ has been chosen to obtain an insignificant barrier height so that the field rolls smoothly to its minimum. A smooth transition can be only obtained for $\lambda\sim{g^{\prime}}^{n}$ with $n\lesssim 3$, while higher powers $n>3$ would increase the barrier height. For $n<3$, the parameter space becomes more restricted by the following constraints: i) to avoid a second period of inflation, and ii) to forbid $\Phi\rightarrow Z^{\prime}Z^{\prime}$ and $\Phi\rightarrow f\bar{f}V$ decays. Therefore, while the assumption of $\lambda\sim{g^{\prime}}^{3}$ represents the best choice of the model, we expect that all the model requirements can be satisfied with a slightly different hierarchy. Nonetheless, assuming a different hierarchy between $\lambda$ and ${g^{\prime}}$ would result in either obtaining a similar parameter space with different numerical values of $g^{\prime}$ or merely in some quantitative changes. The qualitative features in the GW spectral shape would remain unchanged because they are mostly controlled by the $f_{N}$ coupling.